OpenStudy (anonymous):

Determine this quotient. (-2.8) ÷ 4

OpenStudy (alexandervonhumboldt2):

wait, you already posted it?

OpenStudy (anonymous):

yea, this is a new question

OpenStudy (igigighjkl):

I will try to help you> (I'm New Hear).

OpenStudy (anonymous):

thanks

OpenStudy (alexandervonhumboldt2):

\[-2.8/4=-2.8*\frac{ 1 }{ 4 }=-\frac{ 28 }{ 10 }*\frac{ 1 }{ 4 }=\frac{ -28 }{ 40 }=-\frac{ 4*7 }{ 4*10 }=-\frac{ \cancel 4*7 }{ \cancel 4*10}=?\]

OpenStudy (anonymous):

am i supposed to answer that or is it already done...

OpenStudy (alexandervonhumboldt2):

no you have to simplify it

OpenStudy (anonymous):

ohh ok

OpenStudy (igigighjkl):

It equals 0.7

OpenStudy (alexandervonhumboldt2):

cancel 4 and get the answer

OpenStudy (madiimay4music):

YOU FORGOT THE NEGATIIIVEEEE!!!!!!

OpenStudy (igigighjkl):

Did it help you???

OpenStudy (anonymous):

your sure its positive and not negative?

OpenStudy (alexandervonhumboldt2):

it is negative

OpenStudy (madiimay4music):

no it's negatiiiiivvveeee

OpenStudy (igigighjkl):

Oh ok.

OpenStudy (anonymous):

ok ok ok!! thanks!

OpenStudy (anonymous):

okay i will post a new question, it seems more challenging